# DriveMind: A Dual-VLM based Reinforcement Learning Framework for Autonomous Driving

### v1

@@ C.1.2 Vision Language Model Configurations
DriveMind relies on two distinct VLMs: a frozen, high-capacity CLIP model for static semantic anchoring and a compact encoder-decoder model for on-demand prompt generation.
@@
For stable, per-step embedding of bird’s-eye-view observations, we use the OpenCLIP ViT-bigG-14 model. Its image encoder partitions each input frame (resized to $224\times 224$px) into non-overlapping $14\times 14$px patches, projects each patch into a 1 024-dimensional token, and processes the resulting sequence through a 24-layer Vision Transformer. The final CLS token yields a $k=2{,}048$-dimensional embedding
@@
each $\in\mathbb{R}^{2048}$. All weights of ViT-bigG-14 (image and text branches) remain frozen during DriveMind training and evaluation, ensuring a consistent semantic reference frame.
@@
To generate context-specific “present” and “ideal” prompts only when needed, we employ SmolVLM-256M, a lightweight transformer with approximately 256 million parameters. SmolVLM’s vision encoder mirrors a ViT-Tiny architecture (patch size $16\times 16$, input resolution $128\times 128$ px), producing a $d=512$-dimensional latent $z_{t}$. Its autoregressive text decoder is a 6-layer Transformer that, when triggered, generates:
@@
where $P_{t}$ and $I_{t}$ are the natural-language “present” and “ideal” prompts, and $\mathrm{CoT}_{t}$ is the chain-of-thought rationale. These are re-embedded via the SmolVLM text encoder into
@@
and cached until the novelty detector (Eq. A56) signals a scene change. SmolVLM is pretrained via chain-of-thought knowledge distillation from a GPT-4 teacher, minimizing a combined cross-entropy and KL loss that aligns both the prompt outputs and their intermediate rationales.
Text-Prompt Cache Behavior.
A fixed-size cache $\{(u^{P}_{t_{i}},u^{I}_{t_{i}})\}_{i=1}^{K}$ stores the last $K$ dynamic embeddings. Upon each SmolVLM invocation, the oldest entry is evicted and replaced, ensuring that subsequent timesteps reuse these context-specific prompts without redundant decoding. This design keeps real-time inference costs low while retaining up-to-date, scene-adaptive semantic objectives.
@@ C.2 Evaluation Metrics
  • Average Speed (AS): the time-averaged vehicle speed over an episode of length $T$,
@@
    where $v_{\rm col}^{(j)}$ is the vehicle’s speed at the $j$-th collision and $N_{\rm col}$ the total number of collisions.
@@ C.3 Baselines
To provide a comprehensive evaluation of DriveMind, we compare against fourteen state-of-the-art reward methods, organized into three categories. For each, we implement both SAC and PPO where applicable, holding the underlying network architectures and training hyperparameters identical to DriveMind.
C.3.1 Expert-Designed Reward Methods
TIRL-SAC.
TIRL-SAC [26] represents the simplest possible reward: a binary penalty for collisions. At each step $t$, the agent receives
@@ C.3.1 Expert-Designed Reward Methods
This design provides no positive reinforcement for progress, forcing the agent to discover safe behaviors purely by avoiding negative signals. In practice, this extreme sparsity leads to very slow or degenerate learning: without any incentive to move forward, the optimal strategy is often to stay stationary. We implement TIRL-SAC using the standard SAC update rules, with the critic trained to minimize the soft Bellman residual against the sparse $R_{\rm TIRL}$, and the actor updated to maximize a combination of Q-values and policy entropy. All other hyperparameters (learning rate, replay buffer size, batch size, target update rate) match those used for DriveMind.
TIRL-PPO.
TIRL-PPO [26] substitutes the PPO algorithm in place of SAC, but retains the identical binary reward:
@@
where $r_{t}(\phi)=\pi_{\phi}(a_{t}|s_{t})/\pi_{\phi_{\rm old}}(a_{t}|s_{t})$ and $\hat{A}_{t}$ is the estimated advantage using $R_{\rm TIRL\text{-}PPO}$. As with TIRL-SAC, the absence of any positive driving reward causes PPO to converge to near-zero speed policies, highlighting the limitations of purely negative feedback.
Chen-SAC.
Chen-SAC [27] employs a richer, hand-designed reward that balances collision avoidance with smooth, forward driving and lane-keeping:
[…]
Here $v_{t}$ is the vehicle speed, $\Delta_{\rm lane}$ the lateral offset from centerline, and $\omega_{t}$ the steering angle. Typical weights are set to $(w_{\rm coll}=10,\ w_{\rm speed}=1,\ w_{\rm lane}=2,\ w_{\rm steer}=0.5)$. The collision indicator $\mathbf{1}_{\rm coll}$ carries the heaviest penalty to enforce safety. The SAC critic and actor updates proceed as usual, but driven by this composite signal: positive reward for forward motion, negative for deviations and steering effort. In practice, Chen-SAC learns nonzero speeds yet often oscillates in lane-centering, illustrating the difficulty of balancing multiple hand-tuned terms.
ASAP-RL-PPO.
ASAP-RL-PPO [15] focuses on progress and overtaking incentives under PPO:
[…]
Here $\Delta d_{t}$ is the distance advanced along the planned route, $\mathbf{1}_{\rm overtake}$ flags successful passes of slower traffic, and the two penalty indicators discourage collisions and leaving the drivable area. Weight coefficients are chosen as $(\gamma_{\rm prog}=1,\ \gamma_{\rm overtake}=2,\ \gamma_{\rm coll}=5,\ \gamma_{\rm offroad}=3)$. PPO’s clipped surrogate objective encourages stable policy improvements based on these incentives. ASAP-RL-PPO tends to favor higher speeds and overtakes but can be overly aggressive, leading to elevated collision rates.
ChatScene-SAC.
ChatScene-SAC [28] constructs a smoothness-focused reward:
[…]
@@
ChatScene-PPO.
The same smoothness-oriented reward $R_{\rm CS}$ is paired with PPO. PPO’s trajectory batching and clipping help stabilize training but do not fully overcome the conservative bias inherent in the reward design.
C.3.2 LLM-Designed Reward Methods
@@
Revolve [29] leverages a Large Language Model to generate candidate reward functions in Python. The process alternates between prompting the LLM for new code snippets $f_{\theta}$, evaluating them on a small set of human-annotated driving trajectories, and selecting the highest-scoring functions via an evolutionary algorithm. The resulting reward $R_{\rm Revolve}(s_{t},a_{t})=f_{\theta^{*}}(s_{t},a_{t})$ can include complex conditionals, time-based terms, or logical combinations of sensors. When used with SAC or PPO, Revolve often produces unconventional but effective reward structures; however, its reliance on automated code search can introduce brittleness if the LLM generates invalid or overfitted snippets.
Revolve-auto.
Revolve-auto [29] extends Revolve by automatically labeling its own trajectories for subsequent reward refinement. After the initial code is selected, the LLM is re-prompted with performance summaries and asked to generate refinements to $f_{\theta}$. This loop continues for several rounds, yielding $f_{\theta^{\prime}}$ that adapts to the agent’s evolving behavior. While this self-training can improve alignment with driving objectives, it risks reinforcing unintended biases if the feedback loop is not carefully constrained.
C.3.3 VLM-Designed Reward Methods
VLM-SR.
VLM-SR [7] uses the CLIP model to compute cosine similarity between the current image embedding $f_{I}(x_{t})$ and a fixed “goal” text embedding $f_{L}(y^{*})$. Formally,
@@ C.3.3 VLM-Designed Reward Methods
where $\{y_{j}\}$ includes negative distractor prompts and $\tau$ is a temperature. This yields a dense but purely semantic reward, encouraging the agent to produce images “close” to the language goal. In practice, VLM-SR alone struggles in continuous driving, as semantic alignment does not directly capture kinematic feasibility.
@@
RoboCLIP [9] extends VLM-SR to video by averaging per-frame similarities over each episode and penalizing deviations from a task descriptor:
@@
This end-of-episode signal can guide off-policy RL but provides no intermediate shaping, leading to slow convergence in driving tasks.
VLM-RM.
VLM-RM [8] formulates reward as the projection of the current embedding $v_{t}$ onto the vector from a baseline text $u_{\rm base}$ to a target text $u_{\rm target}$:
@@
By measuring progress along the semantic direction, VLM-RM yields a continuous reward. However, without negative goals, it may reward ambiguous or unsafe states that partially match the target description.
@@
LORD-Speed.
LORD-Speed [19] augments LORD with a continuous speed incentive:
@@
VLM-RL.
VLM-RL [20] combines the contrasting language-goal paradigm (positive and negative prompts) with SAC/RL to produce semantic rewards. VLM-RL freezes its VLM, encodes both “present” and “ideal” prompts, and computes
@@ C.3.4 Driving Environments and Scenarios
We evaluate DriveMind in two complementary domains: (i) simulated towns in CARLA for controlled, repeatable benchmarks, and (ii) real-world street scenes drawn from the BDD100K dataset to assess domain transfer.
@@
Let $\mathcal{T}=\{\mathrm{Town}_{1},\dots,\mathrm{Town}_{5}\}$ denote the set of CARLA maps used. During training, the agent is exposed exclusively to $\mathrm{Town}_{2}$, characterized by a compact European-style layout of residential blocks, commercial zones, single-lane roads, and signalized intersections. To formally define each episode’s environment, we sample a scenario by:
@@
Real-World BDD100K Scenes.
To quantify domain transfer, we collect a set of $N=10{,}000$ bird’s-eye-view (BEV) projections from the BDD100K validation split. Let
=$\mathcal{I}=\{\,I_{i}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\dots,N\}$ (A49)
be the set of images, each paired with ground-truth lane and vehicle masks. We define an evaluation scenario for each $I_{i}$ by treating it as a static scene with $D=\text{empty}$ and no navigation route. The policy executes a single control step $a_{i}$ from its observation $o_{i}=(I_{i},e_{i},\varnothing)$, and we record the positive and negative VLM similarity scores.
@@
During both simulated and real-world evaluation, scenarios are drawn uniformly from their respective sets $\mathcal{T}\times\mathcal{D}\times\mathcal{R}$ or $\mathcal{I}$. Performance is reported as mean $\pm$ standard deviation over three random seeds (simulated) or three disjoint image subsets (real). In CARLA, we compute driving efficiency metrics (average speed, route completion, distance traveled) and safety metrics (collision rate, collision speed), while in BDD100K we report the mean positive/negative score and the distributional distances $(W,D_{\mathrm{KS}})$ to quantify domain gap.
@@ C.3.5 Navigation and Novelty Triggering
where $L(R)$ is the Euclidean length of the polyline $R=(R_{1},\dots,R_{|R|})$. If $L(R)$ exceeds a pre-specified maximum $L_{\max}$ (e.g. 3 km), we reject and resample $(p_{\mathrm{start}},p_{\mathrm{goal}})$ until $L(R)\leq L_{\max}$. By continuously re-sampling upon completion or termination, the agent experiences diverse navigation challenges within a bounded episode length.
@@
and maintain a cache =$\mathcal{B}=\{v^{(1)},\ldots,v^{(K)}\}$ of the $K$ most recent embeddings. We define the novelty score
@@
if $\text{if }\Delta_{t}>\delta,\quad\text{then invoke SmolVLM on }\psi(s_{t})\text{ to generate new prompts;}$; (A57)
otherwise, reuse the last cached dynamic embeddings $(u^{P}_{t},u^{I}_{t})$. Upon invocation, we update the cache via a sliding-window policy:
($\mathcal{B}\;\leftarrow\;\bigl{(}\mathcal{B}\setminus\{v^{(1)}\}\bigr{)}\;\cup\;\{v_{t}\},$ (A58)
dropping the oldest embedding and appending the current one. In practice, $\delta$ is chosen as a high quantile of the distribution of $\Delta_{t}$ observed during a warm-up phase, ensuring that SmolVLM is called only for genuinely unfamiliar or rare scenes. This mechanism keeps the per-step complexity low while granting DriveMind the ability to adapt its semantic objectives precisely when the driving context changes.

### v2

@@ C.1.2 Vision Language Model Configurations
DriveMind relies on two distinct VLMs: a frozen, high‑capacity CLIP model for static semantic anchoring and a compact encoder-decoder model for on‑demand prompt generation.
@@
For stable, per‑step embedding of bird’s‑eye‑view observations, we use the OpenCLIP ViT‑bigG‑14 model. Its image encoder partitions each input frame (resized to $224\times 224$px) into non‑overlapping $14\times 14$px patches, projects each patch into a 1 024‑dimensional token, and processes the resulting sequence through a 24‑layer Vision Transformer. The final CLS token yields a k=2,048k=2{,}048‑dimensional embedding
@@
each $\in\mathbb{R}^{2048}$. All weights of ViT‑bigG‑14 (image and text branches) remain frozen during DriveMind training and evaluation, ensuring a consistent semantic reference frame.
@@
To generate context‑specific “present” and “ideal” prompts only when needed, we employ SmolVLM‑256M, a lightweight transformer with approximately 256 million parameters. SmolVLM’s vision encoder mirrors a ViT‑Tiny architecture (patch size $16\times 16$, input resolution $128\times 128$ px), producing a d=512d=512‑dimensional latent $z_{t}$. Its autoregressive text decoder is a 6‑layer Transformer that, when triggered, generates:
@@
where $P_{t}$ and $I_{t}$ are the natural‑language “present” and “ideal” prompts, and $\mathrm{CoT}_{t}$ is the chain‑of‑thought rationale. These are re‑embedded via the SmolVLM text encoder into
@@
and cached until the novelty detector (Eq. A56) signals a scene change. SmolVLM is pretrained via chain‑of‑thought knowledge distillation from a GPT-4 teacher, minimizing a combined cross‑entropy and KL loss that aligns both the prompt outputs and their intermediate rationales.
Text‑Prompt Cache Behavior.
A fixed‐size cache $\{(u^{P}_{t_{i}},u^{I}_{t_{i}})\}_{i=1}^{K}$ stores the last $K$ dynamic embeddings. Upon each SmolVLM invocation, the oldest entry is evicted and replaced, ensuring that subsequent timesteps reuse these context‑specific prompts without redundant decoding. This design keeps real‑time inference costs low while retaining up‑to‑date, scene‑adaptive semantic objectives.
@@ C.2 Evaluation Metrics
  • Average Speed (AS): the time‑averaged vehicle speed over an episode of length $T$,
@@
    where $v_{\rm col}^{(j)}$ is the vehicle’s speed at the jj‑th collision and $N_{\rm col}$ the total number of collisions.
@@ C.3 Baselines
To provide a comprehensive evaluation of DriveMind, we compare against fourteen state‑of‑the‑art reward methods, organized into three categories. For each, we implement both SAC and PPO where applicable, holding the underlying network architectures and training hyperparameters identical to DriveMind.
C.3.1 Expert‑Designed Reward Methods
TIRL‑SAC.
TIRL‑SAC [2] represents the simplest possible reward: a binary penalty for collisions. At each step $t$, the agent receives
@@ C.3.1 Expert‑Designed Reward Methods
This design provides no positive reinforcement for progress, forcing the agent to discover safe behaviors purely by avoiding negative signals. In practice, this extreme sparsity leads to very slow or degenerate learning: without any incentive to move forward, the optimal strategy is often to stay stationary. We implement TIRL‑SAC using the standard SAC update rules, with the critic trained to minimize the soft Bellman residual against the sparse $R_{\rm TIRL}$, and the actor updated to maximize a combination of Q‑values and policy entropy. All other hyperparameters (learning rate, replay buffer size, batch size, target update rate) match those used for DriveMind.
TIRL‑PPO.
TIRL‑PPO [2] substitutes the PPO algorithm in place of SAC, but retains the identical binary reward:
@@
where $r_{t}(\phi)=\pi_{\phi}(a_{t}|s_{t})/\pi_{\phi_{\rm old}}(a_{t}|s_{t})$ and $\hat{A}_{t}$ is the estimated advantage using $R_{\rm TIRL\text{-}PPO}$. As with TIRL‑SAC, the absence of any positive driving reward causes PPO to converge to near‐zero speed policies, highlighting the limitations of purely negative feedback.
Chen‑SAC.
Chen‑SAC [4] employs a richer, hand‑designed reward that balances collision avoidance with smooth, forward driving and lane‐keeping:
$R_{\rm Chen}(s_{t},a_{t})=-\,w_{\rm coll}\,\mathbf{1}_{\rm coll}(s_{t})+w_{\rm speed}\,\frac{v_{t}}{v_{\max}}-w_{\rm lane}\,\Delta_{\rm lane}(s_{t})-w_{\rm steer}\,\lvert\omega_{t}\rvert$. (A38)
Here $v_{t}$ is the vehicle speed, $\Delta_{\rm lane}$ the lateral offset from centerline, and $\omega_{t}$ the steering angle. Typical weights are set to $(w_{\rm coll}=10,\ w_{\rm speed}=1,\ w_{\rm lane}=2,\ w_{\rm steer}=0.5)$. The collision indicator $\mathbf{1}_{\rm coll}$ carries the heaviest penalty to enforce safety. The SAC critic and actor updates proceed as usual, but driven by this composite signal: positive reward for forward motion, negative for deviations and steering effort. In practice, Chen‑SAC learns nonzero speeds yet often oscillates in lane‐centering, illustrating the difficulty of balancing multiple hand‑tuned terms.
ASAP‑RL‑PPO.
ASAP‑RL‑PPO [22] focuses on progress and overtaking incentives under PPO:
$R_{\rm ASAP}(s_{t},a_{t})=\gamma_{\rm prog}\,\Delta d_{t}+\gamma_{\rm overtake}\,\mathbf{1}_{\rm overtake}(s_{t})-\gamma_{\rm coll}\,\mathbf{1}_{\rm coll}(s_{t})-\gamma_{\rm offroad}\,\mathbf{1}_{\rm offroad}(s_{t}).$ (A39)
Here $\Delta d_{t}$ is the distance advanced along the planned route, $\mathbf{1}_{\rm overtake}$ flags successful passes of slower traffic, and the two penalty indicators discourage collisions and leaving the drivable area. Weight coefficients are chosen as $(\gamma_{\rm prog}=1,\ \gamma_{\rm overtake}=2,\ \gamma_{\rm coll}=5,\ \gamma_{\rm offroad}=3)$. PPO’s clipped surrogate objective encourages stable policy improvements based on these incentives. ASAP‑RL‑PPO tends to favor higher speeds and overtakes but can be overly aggressive, leading to elevated collision rates.
ChatScene‑SAC.
ChatScene‑SAC [27] constructs a smoothness‐focused reward:
$R_{\rm CS}(s_{t},a_{t})=-\,\lambda_{1}\bigl|\ddot{x}_{t}\bigr|-\,\lambda_{2}\lvert a_{y,t}\rvert-\,\lambda_{3}\,\mathbf{1}_{\rm coll}(s_{t})+\epsilon,$ (A40)
@@
ChatScene‑PPO.
The same smoothness‐oriented reward $R_{\rm CS}$ is paired with PPO. PPO’s trajectory batching and clipping help stabilize training but do not fully overcome the conservative bias inherent in the reward design.
C.3.2 LLM‑Designed Reward Methods
@@
Revolve [9] leverages a Large Language Model to generate candidate reward functions in Python. The process alternates between prompting the LLM for new code snippets $f_{\theta}$, evaluating them on a small set of human‐annotated driving trajectories, and selecting the highest‐scoring functions via an evolutionary algorithm. The resulting reward $R_{\rm Revolve}(s_{t},a_{t})=f_{\theta^{*}}(s_{t},a_{t})$ can include complex conditionals, time‐based terms, or logical combinations of sensors. When used with SAC or PPO, Revolve often produces unconventional but effective reward structures; however, its reliance on automated code search can introduce brittleness if the LLM generates invalid or overfitted snippets.
Revolve‑auto.
Revolve‑auto [9] extends Revolve by automatically labeling its own trajectories for subsequent reward refinement. After the initial code is selected, the LLM is re‑prompted with performance summaries and asked to generate refinements to $f_{\theta}$. This loop continues for several rounds, yielding $f_{\theta^{\prime}}$ that adapts to the agent’s evolving behavior. While this self‑training can improve alignment with driving objectives, it risks reinforcing unintended biases if the feedback loop is not carefully constrained.
C.3.3 VLM‑Designed Reward Methods
VLM‑SR.
VLM‑SR [1] uses the CLIP model to compute cosine similarity between the current image embedding $f_{I}(x_{t})$ and a fixed “goal” text embedding $f_{L}(y^{*})$. Formally,
@@ C.3.3 VLM‑Designed Reward Methods
where $\{y_{j}\}$ includes negative distractor prompts and $\tau$ is a temperature. This yields a dense but purely semantic reward, encouraging the agent to produce images “close” to the language goal. In practice, VLM‑SR alone struggles in continuous driving, as semantic alignment does not directly capture kinematic feasibility.
@@
RoboCLIP [19] extends VLM‑SR to video by averaging per‐frame similarities over each episode and penalizing deviations from a task descriptor:
@@
This end‐of‐episode signal can guide off‐policy RL but provides no intermediate shaping, leading to slow convergence in driving tasks.
VLM‑RM.
VLM‑RM [16] formulates reward as the projection of the current embedding $v_{t}$ onto the vector from a baseline text $u_{\rm base}$ to a target text $u_{\rm target}$:
@@
By measuring progress along the semantic direction, VLM‑RM yields a continuous reward. However, without negative goals, it may reward ambiguous or unsafe states that partially match the target description.
@@
LORD‑Speed.
LORD‑Speed [26] augments LORD with a continuous speed incentive:
@@
VLM‑RL.
VLM‑RL [11] combines the contrasting language‑goal paradigm (positive and negative prompts) with SAC/RL to produce semantic rewards. VLM‑RL freezes its VLM, encodes both “present” and “ideal” prompts, and computes
@@ C.3.4 Driving Environments and Scenarios
We evaluate DriveMind in two complementary domains: (i) simulated towns in CARLA for controlled, repeatable benchmarks, and (ii) real‐world street scenes drawn from the BDD100K dataset to assess domain transfer.
@@
Let $\mathcal{T}=\{\mathrm{Town}_{1},\dots,\mathrm{Town}_{5}\}$ denote the set of CARLA maps used. During training, the agent is exposed exclusively to $\mathrm{Town}_{2}$, characterized by a compact European‐style layout of residential blocks, commercial zones, single‐lane roads, and signalized intersections. To formally define each episode’s environment, we sample a scenario by:
@@
Real‐World BDD100K Scenes.
To quantify domain transfer, we collect a set of N=10, [^0] N=10{,}000 bird’s‐eye‐view (BEV) projections from the BDD100K validation split. Let
$\mathcal{I}=\{\,I_{i}\in\mathbb{R}^{H\times W\times 3}\mid i=1,\dots,N\}$ (A49)
be the set of images, each paired with ground‐truth lane and vehicle masks. We define an evaluation scenario for each $I_{i}$ by treating it as a static scene with $D=\text{empty}$ and no navigation route. The policy executes a single control step $a_{i}$ from its observation $o_{i}=(I_{i},e_{i},\varnothing)$, and we record the positive and negative VLM similarity scores.
@@
During both simulated and real‐world evaluation, scenarios are drawn uniformly from their respective sets $\mathcal{T}\times\mathcal{D}\times\mathcal{R}$ or $\mathcal{I}$. Performance is reported as mean $\pm$ standard deviation over three random seeds (simulated) or three disjoint image subsets (real). In CARLA, we compute driving efficiency metrics (average speed, route completion, distance traveled) and safety metrics (collision rate, collision speed), while in BDD100K we report the mean positive/negative score and the distributional distances $(W,D_{\mathrm{KS}})$ to quantify domain gap.
@@ C.3.5 Navigation and Novelty Triggering
where $L(R)$ is the Euclidean length of the polyline $R=(R_{1},\dots,R_{|R|})$. If $L(R)$ exceeds a pre‐specified maximum $L_{\max}$ (e.g. 3 km), we reject and resample $(p_{\mathrm{start}},p_{\mathrm{goal}})$ until $L(R)\leq L_{\max}$. By continuously re‐sampling upon completion or termination, the agent experiences diverse navigation challenges within a bounded episode length.
@@
and maintain a cache $\mathcal{B}=\{v^{(1)},\ldots,v^{(K)}\}$ of the $K$ most recent embeddings. We define the novelty score
@@
[…]
otherwise, reuse the last cached dynamic embeddings $(u^{P}_{t},u^{I}_{t})$. Upon invocation, we update the cache via a sliding‐window policy:
$\mathcal{B}\;\leftarrow\;\bigl(\mathcal{B}\setminus\{v^{(1)}\}\bigr)\;\cup\;\{v_{t}\},$ (A58)
dropping the oldest embedding and appending the current one. In practice, $\delta$ is chosen as a high quantile of the distribution of $\Delta_{t}$ observed during a warm‐up phase, ensuring that SmolVLM is called only for genuinely unfamiliar or rare scenes. This mechanism keeps the per‐step complexity low while granting DriveMind the ability to adapt its semantic objectives precisely when the driving context changes.